OpenStudy (anonymous):

4x^-2y^-4z^-3 divided by 12x^-5y^-4z^2

OpenStudy (anonymous):

^ is the exponent sign. Can you re-type it a bit clearer with more brackets?

OpenStudy (anonymous):

@jama333 do you know how to solve this system of equations using elimination -4x-2y=12 and 4x+8y=-24

OpenStudy (anonymous):

the choices a.(5,3) b.(-6,-6) c.(3,1) d.(6,-6)

OpenStudy (anonymous):

First you simplify each equation. The first one can be divided by 2 to get -2x-y=6 The second one can be divided by 4 to get x+2y=-6

OpenStudy (anonymous):

^ means there powers, what about those?

OpenStudy (anonymous):

\[4x ^{-2}y ^{4}z ^{-3} \div 12x ^{-5}y ^{-4}z ^{2}\]

OpenStudy (anonymous):

I solved the two equations, but I get (-2,-2) which isn't one of your choices. Are you sure you typed the two equations correctly. You typed. -4x-2y=12 and 4x+8y=-24 As for the other comment, yes this (^) symbol means powers of exponents. When you typed it in the initial post, you typed: 4x^-2y^-4z^-3 divided by 12x^-5y^-4z^2 If all of the exponents are negative except the z^2, then I know what to do. I just wanted to make sure you typed it right.

OpenStudy (anonymous):

I just corrected it..if that makes it easier

OpenStudy (anonymous):

Okay so for the first term, leave the 4y^4 in the numerator and put everything else in the denominator. When you put it in the denominator, it is no longer a negative exponent. So the first term becomes (4y^4) / [(x^2)(z^3)]

OpenStudy (anonymous):

The second term becomes (12z^2) / [(x^5)(y^4)].

OpenStudy (anonymous):

So remember that x^(-2) = 1/(x^2) So now that we've gotten rid of the negative exponents, it's easier to deal with it. When you multiply by a fraction, it's the same as dividing by its reciprocal. The reciprocal of 2/3 is 3/2.

OpenStudy (anonymous):

(4y^4) / (x^2)(z^3) TIMES (x^5)(y^4) / 12z^2 Do you know you to cancel stuff?

OpenStudy (anonymous):